The reading speed of second grade students in a large city is approximately​ normal, with a mean of 89 words per minute​ (wpm) and a standard deviation of 10 wpm. Complete parts​ (a) through​ (e).
​(a) What is the probability a randomly selected student in the city will read more than 93 words per​ minute?
The probability is
0.3446
​(Round to four decimal places as​ needed.)
​(b) What is the probability that a random sample of 13 second grade students from the city results in a mean reading rate of more than 93 words per​ minute?

Answer by stanbon(75887)

The reading speed of second grade students in a large city is approximately​ normal, with a mean of 89 words per minute​ (wpm) and a standard deviation of 10 wpm. Complete parts​ (a) through​ (e).
​(a) What is the probability a randomly selected student in the city will read more than 93 words per​ minute?
The probability is
0.3446
​(Round to four decimal places as​ needed.)
0.3446 is correct
z(93) = (93-89)/10 = 0.4
Ans:: P(x > 93) = P(z > 0.4) = normalcdf(0.4,100) = 0.3664
---------------------------------------------
​(b) What is the probability that a random sample of 13 second grade from the city results in a mean reading rate of more than 93 words per​ minute?
z(93) = (93-89)/(10/sqrt(13)) = sqrt(13)*0.4 = 1.4422
Ans: p(x-bar > 93) = P(z > 1.4422) = normalcdf(1.4422,100) = 0.0746
-----
Cheers,
Stan H.
----------